For the last

3

years

Kerwin

has made deposits of

$120.00

at the end of every

month

earning interest at

10%

compounded

monthly.

If he leaves the accumulated balance for another

5

years at

9%

compounded

quarterly

what will the balance be in the account?
Answer #1

PMT =120
Rate per month =10%/12
Number of Periods =12*5 =60
FV using 60 periods =PMT*(((1+r)%n-1)/r) =120*(((1+10%/12)^60-1)/(10%/12))=9292.4487

Rate per quarter =(9%/4) =2.25%
Number of Periods =5*4 =20
FV after 10 years =9292.4487*(1+2.25%)^20 =14500.95